OpenStudy (anonymous):

If it takes 3 days for 10 workers to finish building one house, how may days will it take 15 workers to finish 4 houses?

OpenStudy (anonymous):

i hate word problems

OpenStudy (anonymous):

me too =(

OpenStudy (anonymous):

we can find the unit rate per worker i guess

OpenStudy (anonymous):

hows that? all i know is, is that they build 1 house in 3 days so 1/3

OpenStudy (anonymous):

ten workers can build one third of a house per day

OpenStudy (anonymous):

so each worker builds \(\frac{1}{30}\) of a house per day that is the unit rate per worker

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

how did u get that/. and y do u divide or multiply?

OpenStudy (anonymous):

i divided \(\frac{1}{3}\) by \(10\) to get \(\frac{1}{30}\)

OpenStudy (anonymous):

but why do u divide instead of multiplying?

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

k

OpenStudy (anonymous):

you already said the \(\frac{1}{3}\) of a house per day right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but that was not for one worker, that is for 10 workers working together

OpenStudy (anonymous):

ohhhhhhh i see so to find how much they get done PER WORKER we have to divide!!

OpenStudy (anonymous):

so each single worker does one -tenth of that work i.e. \[\frac{1}{3}\times \frac{1}{10}=\frac{1}{30}\]

OpenStudy (anonymous):

right divide by 10 or multiply by 1/10 same thing

OpenStudy (anonymous):

we are still not done though

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

now you have 15 workers right?

OpenStudy (anonymous):

yea so multiply by 15 to see how much work 15 workers can do

OpenStudy (anonymous):

exactly \[15\times \frac{1}{30}=?\]

OpenStudy (anonymous):

2/15

OpenStudy (anonymous):

hmm no careful

OpenStudy (anonymous):

what is fifteen times one thirtieth?

OpenStudy (anonymous):

i simpilfied

OpenStudy (anonymous):

not correctly though

OpenStudy (anonymous):

15/30= 1=2

OpenStudy (anonymous):

oops 1/2

OpenStudy (anonymous):

1/2 right

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

they build half a house per day one how in two days

OpenStudy (anonymous):

how long before they build 4 houses?

OpenStudy (anonymous):

so do 1/2 * 4?

OpenStudy (anonymous):

again lets go slow

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

they build half a house per day right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so one house takes how long to build?

OpenStudy (anonymous):

1 house every 2 days

OpenStudy (anonymous):

right

OpenStudy (anonymous):

if it takes two days to build one house, how long to build four?

OpenStudy (anonymous):

proportions

OpenStudy (anonymous):

?

OpenStudy (anonymous):